Pattern Problem
Determine the next number in the sequence 0, 10, 24, 56, 112, 190, ...

Answer by math_helper(2461)


Sequence 0, 10, 24, 56, 112, 190, ...
first differences: 10 14 32 56 78
2nd differences: 4 18 24 22
3rd differences: 14 6 -2
4th differences: -8 -8
So the next 3rd diff is -2-8 = -10
The next 2nd diff is 22-10 = 12
The next first difference 78+12 = 90
The next number in sequence is 190 + 90 = 280